# Positional encoding is not the same as context: A study on positional encoding for Sequential recommendation

### v1

@@ A.1 Experiments
Table 10: Games Dataset results with different positional encodings. Original results from Rashed et al. (2022) at the end.
A.2 Resources
Due to the instability of these models, we had to run many seeds to get some confident results. We started by running each one of the model configurations three times. After that, we decided to run more seeds based on the confidence of each one of the cases. We tried to keep at least a confidence range of ten points. We used $V100$ GPUs of 32GBs. For $men$ and $fashion$, based on the encoding, the training times go from $4$ to $6$ hours. For $Beauty$, they went from $22$ to $26$ hours. Based on the team’s resource ability, these experiments were carried out over $6$ months, lasting between $1$ and $8$.
We used the following seeds: $42$, $43$, $44$, $45$, $46$, $1809$, $1810$, $1811$, $1812$, $1741$, $1742$, $1743$, $1744$, $1745$, $123$, $234$, $345$, $456$, $567$, $678$, $789$, $890,102$, $203$, $304$, $405$, $506$, $607$, $708$, $808$, $901$, $1001$, $987$, $8765$, $7654$, and $6543$.
A.3 Hyperparameters
In this section, we present the hyperparameters used for training. As mentioned above, we use the same hyperparameters as CARCA Rashed et al. (2022). In some cases, we experimented with the upper bound from the norm in the encoding and the training length.
@@ A.3 Hyperparameters
Table 11: Hyperparameters with corresponding options were used in the original CARCA model. This parameter were written in tensor flow. In "L2 reg. weight", the value 00 corresponds to no bound, which in Pytorch is $Nan$, and we denote it as None .
A.4 Deviations over all the trainings
Here, we present the table with all the deviation values from the training. In the table in the main text, we considered those training where we use the default values for the upper bound, i.e., $0.0001$, and $None$ for the Games dataset. We can see that the trends pointed out in Section 5.1 are still valid in this table. The encodings with concatenation are more stable, and RMHA-4 is the most stable one. We can see that with longer trainings Rotatory encodings become more stable than Rope
encoding HIT dev NDCG dev runs CI-length
@@ A.4 Deviations over all the trainings
Table 12: Deviations among the different encodings. Those per head, like RMHA-4 or Rope , are more stable than others.
A.5 Losses
@@
Figure 2: Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ ad silu. Dataset: Fashion
Figure 3: Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ ad silu. Dataset: Men

### v2

@@ A.1. Experiments
Table 12. Games Dataset results with different positional encodings. Results from (Rashed et al., 2022) at the end.
A.2. Sparsity and deviation experiments
In this section, we present tables of the datasets created by reduction to evaluate whether the encoding type depends on the vector representation of the items, deviation, and sparsity. For these cases, we considered only three encodings: RMHA-4 , Rotatory , and None . Table 17 contains results for datasets with a deviation below $12$, which we deemed invalid and have therefore moved to this appendix.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
leaky None 0.0001 48.00 22.19 33.53 26.31 6 (30.24, 65.76) 35.52
silu None 0.0001 66.30 18.61 55.17 21.99 6 (51.41, 81.19) 29.78
leaky None 0.1000 32.94 1.89 16.77 3.00 6 (31.43, 34.45) 3.02
silu None 0.1000 63.85 17.73 52.02 20.44 6 (49.66, 78.04) 28.38
leaky Rotatory + Con 0.0001 50.04 19.62 35.82 22.95 6 (34.34, 65.74) 31.40
silu Rotatory + Con 0.0001 63.87 19.99 52.59 22.77 6 (47.87, 79.87) 32.00
leaky Rotatory + Con 0.1000 49.17 21.02 35.89 25.10 6 (32.35, 65.99) 33.64
silu Rotatory + Con 0.1000 67.94 19.26 57.24 23.51 6 (52.53, 83.35) 30.82
leaky RMHA-4 0.0001 79.72 2.85 71.57 3.66 6 (77.44, 82.0) 4.56
silu RMHA-4 0.0001 46.55 14.12 31.68 16.59 6 (35.25, 57.85) 22.60
leaky RMHA-4 0.1000 57.99 20.07 45.36 24.01 6 (41.93, 74.05) 32.12
silu RMHA-4 0.1000 61.14 14.64 48.46 18.10 6 (49.43, 72.85) 23.42
Table 13. SubFashion Dataset results for the best positional encodings and the None encoding case.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
leaky None 0.0001 77.10 3.72 67.89 5.25 12 (75.0, 79.2) 4.20
silu None 0.0001 76.33 6.11 67.00 7.08 12 (72.87, 79.79) 6.92
leaky None 0.1000 69.29 15.21 58.64 18.19 6 (57.12, 81.46) 24.34
silu None 0.1000 74.66 4.13 64.38 5.85 6 (71.36, 77.96) 6.60
leaky Rotatory + Con 0.0001 71.82 11.90 61.69 14.11 12 (65.09, 78.55) 13.46
silu Rotatory + Con 0.0001 75.28 3.63 65.31 5.03 12 (73.23, 77.33) 4.10
leaky Rotatory + Con 0.1000 73.93 5.87 63.77 8.00 6 (69.23, 78.63) 9.40
silu Rotatory + Con 0.1000 73.11 5.43 62.64 6.80 6 (68.77, 77.45) 8.68
leaky RMHA-4 0.0001 75.96 2.71 66.03 4.00 8 (74.08, 77.84) 3.76
silu RMHA-4 0.0001 70.44 9.61 59.67 11.50 8 (63.78, 77.1) 13.32
leaky RMHA-4 0.1000 77.28 2.66 68.45 3.65 6 (75.15, 79.41) 4.26
silu RMHA-4 0.1000 72.54 4.57 61.75 5.96 6 (68.88, 76.2) 7.32
Table 14. SubMen Dataset results for the best positional encodings and the None encoding case.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
leaky None NaN 49.53 1.01 28.84 0.69 6 (48.72, 50.34) 1.62
silu None NaN 48.31 1.22 28.36 0.86 6 (47.33, 49.29) 1.96
leaky Rotatory + Con NaN 52.23 1.57 31.51 0.96 6 (50.97, 53.49) 2.52
silu Rotatory + Con NaN 50.73 0.89 30.48 0.72 6 (50.02, 51.44) 1.42
leaky RMHA-4 NaN 50.43 0.95 29.96 0.87 6 (49.67, 51.19) 1.52
silu RMHA-4 NaN 48.97 1.35 28.57 0.79 6 (47.89, 50.05) 2.16
Table 15. Subgames Dataset results for the best positional encodings (Rotatory + Con and RMHA-4 ) and the None encoding case.
Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev runs CI CI-length
leaky None 0.0001 72.08 6.04 61.47 7.96 6 (67.25, 76.91) 9.66
silu None 0.0001 73.95 3.33 64.05 4.71 6 (71.29, 76.61) 5.32
leaky None 0.1000 71.56 5.38 60.96 6.93 6 (67.26, 75.86) 8.60
silu None 0.1000 70.51 6.04 59.29 7.69 6 (65.68, 75.34) 9.66
leaky Rotatory + Con 0.0001 71.52 6.39 60.71 8.17 6 (66.41, 76.63) 10.22
silu Rotatory + Con 0.0001 72.77 2.96 62.56 4.31 6 (70.4, 75.14) 4.74
leaky Rotatory + Con 0.1000 73.12 2.38 62.86 2.99 6 (71.22, 75.02) 3.80
silu Rotatory + Con 0.1000 72.03 3.90 61.33 4.76 6 (68.91, 75.15) 6.24
silu RMHA-4 0.0001 53.67 8.81 36.15 12.73 12 (48.69, 58.65) 9.96
leaky RMHA-4 0.1000 61.82 12.74 48.93 14.82 6 (51.63, 72.01) 20.38
silu RMHA-4 0.1000 61.27 5.87 47.24 6.96 6 (56.57, 65.97) 9.40
Table 16. SubMen2 Dataset results for the best positional encodings and the None encoding case.
Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
Submen leaky RMHA-4 0.1000 77.28 2.66 68.45 3.65
Submen leaky None 0.0001 77.10 3.72 67.89 5.25
Submen silu None 0.0001 76.33 6.11 67.00 7.08
Submen leaky RMHA-4 0.0001 75.96 2.71 66.03 4.00
Subfashion leaky RMHA-4 0.0001 79.72 2.85 71.57 3.66
Subfashion leaky None 0.1000 32.94 1.89 16.77 3.00
Subgames leaky Rotatory + Con NaN 52.23 1.57 31.51 0.96
Subgames silu Rotatory + Con NaN 50.73 0.89 30.48 0.72
Subgames leaky RMHA-4 NaN 50.43 0.95 29.96 0.87
Subgames leaky None NaN 49.53 1.01 28.84 0.69
Submen3 silu None 0.0001 73.95 3.33 64.05 4.71
Submen3 leaky Rotatory + Con 0.1000 73.12 2.38 62.86 2.99
Submen3 silu Rotatory + Con 0.0001 72.77 2.96 62.56 4.31
Submen3 leaky None 0.0001 72.08 6.04 61.47 7.96
Table 17. The table displays the top $4$ results for each dataset on cases with a standard deviation for HIT below $12$.
A.3. Resources
Due to the inherent instability of these models, it was necessary to run multiple seeds to obtain reliable results. Initially, we executed each model configuration three times. However, as we analyzed the results, we determined that additional seeds were needed for cases with high-deviation, aiming to maintain a confidence interval within ten points.
All experiments were conducted on $V100$ GPUs with $32$ GB of memory. Training times varied depending on the dataset and encoding type. For the Men and Fashion datasets, training durations ranged from $4$ to $6$ hours, while for Beauty, the training times extended from $22$ to $26$ hours. Given the available resources, these experiments were spread over a period of six months.
We used the following seed values to ensure diversity in our training runs: $42$, $43$, $44$, $45$, $46$, $1809$, $1810$, $1811$, $1812$, $1741$, $1742$, $1743$, $1744$, $1745$, $123$, $234$, $345$, $456$, $567$, $678$, $789$, $890,102$, $203$, $304$, $405$, $506$, $607$, $708$, $808$, $901$, $1001$, $987$, $8765$, $7654$, and $6543$.
A.4. Hyperparameters
In this section, we outline the hyperparameters used during training. Consistent with the approach used in CARCA (Rashed et al., 2022), we adopted the same baseline hyperparameters across most experiments to maintain comparability. In specific cases, we experimented with variations in the upper bound of the norm for the encodings, activation functions, as well as adjusting the training duration to explore its impact on stability and performance.
@@ A.4. Hyperparameters
Table 18. Hyperparameters with corresponding options were used in the original CARCA model. This parameter were written in tensor flow. In "$L2$ reg. weight", the value 00 corresponds to no bound, which in Pytorch is $Nan$, and we denote it as None .
A.5. Deviations over all the trainings
In this section, we present the complete table of deviation values from all training sessions. In the main text, we focused on the training runs where the default upper bound values ($0.0001$) were used, along with $None$ for the Games dataset. The trends discussed in Section 5.1 remain consistent here. Encodings that incorporate concatenation generally demonstrate greater stability, with RMHA-4 continuing to be the most stable overall. Additionally, we observe that with extended training durations, the Rotatory encodings surpass RoPE in stability.
encoding Avg Dev Hit Avg Dev NDCG runs CI-length
@@ A.5. Deviations over all the trainings
Table 19. Deviations among the different encodings. Those per head, like RMHA-4 or RoPE , are more stable than others, together with Rotatory .
A.6. Losses
@@
Figure 4. Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ and silu. Dataset: Fashion
Figure 5. Losses randomly selected from None , first row, and RMHA-4 , second row, for the test set with $0.0001$ and silu. Dataset: Men
A.7. Decision diagram
Figure 6. Decision tree for encoding selection based on dataset deviation. The diagram outlines the recommended encodings.
A.8. Extra experiments with SASRec++
All of our experiments were conducted using the CARCA model (Rashed et al., 2022). To assess whether the encoding behaviors observed in CARCA extend to other architectures, we also evaluated the SASRec++ model, an extension of the original SASRec model (Kang and McAuley, 2018a). SASRec+++++ + is conceptually similar to CARCA but differs in that it replaces the transformer decoder with a simple dot product operation.
We applied SASRec++ to the Submen dataset using the following encodings: None , Learnt + Con , Rotatory + Con , Abs + Con , Learnt , Abs , RMHA-4 , RopeOne , RoPE , and Rotatory . As shown in Table 20, we find that RMHA-4 continues to outperform other encodings, maintaining its position as the best-performing encoding across all experiments. Most of the results cluster around $40\%$, with minimal differences observed between encodings. The difference between RopeOne and other encodings is particularly small, with only slight improvements seen for RMHA-4 with a low Hit Dev  and Learnt + Con (as seen in Table 21) with a higher Hit Dev.
Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
Submen silu RMHA-4 0.1000 44.12 2.58 19.73 1.17
Submen silu RMHA-4 0.0001 42.45 2.91 18.91 1.11
Submen leaky RopeOne 0.0001 40.69 0.75 18.90 0.60
Submen leaky RopeOne 0.1000 40.66 0.74 18.67 0.39
Table 20. The table displays the top $4$ results for Submen for SASRec++ on cases with a standard deviation for Hit Meanbelow $3$.
Dataset Act encoding nmax Hit Mean Hit Dev NDCG Mean NDCG Dev
Submen silu Learnt + Con 0.0001 58.08 11.49 26.41 6.37
Submen silu RMHA-4 0.1000 44.12 2.58 19.73 1.17
Submen leaky RMHA-4 0.0001 42.86 5.52 20.57 5.34
Submen silu RMHA-4 0.0001 42.45 2.91 18.91 1.11
Table 21. The table displays the top $4$ results for Submen for SASRec++ on cases with a standard deviation for Hit Meanbelow $12$.